what is the maximum value of the function whose derivative posseses no roots?

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Derivative posses no roots => Derivative is constant
\frac{dy}{dx} = constant = c
dy = c.dx
\int dy = \int c.dx
y = cx + d, d = constant
Maximum value is infinite if we are taking the domain to be R.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty